# DropleX: Liquid sensing on tablet touchscreens

### v3

@@ Abstract.
We present DropleX, the first system that enables liquid sensing using the capacitive touchscreen of commodity tablets. DropleX detects microliter–scale liquid samples, and performs non-invasive, through-container measurements to detect whether a drink has been spiked or if a sealed liquid has been contaminated. These capabilities are made possible by a physics-informed mechanism that disables the touchscreen’s built-in adaptive filters, originally designed to reject the effects of liquid drops such as rain, without any hardware modifications. We model the touchscreen’s sensing capabilities, limits, and non-idealities to inform the design of a signal processing and learning-based pipeline for liquid sensing. Our system achieves 96–99% accuracy in detecting microliter-scale adulteration in soda, wine, and milk, 93–96% accuracy in threshold detection of trace chemical concentrations, and 86–96% accuracy in through-container adulterant detection. Given the predominance of touchscreens, these exploratory results can open new opportunities for liquid sensing on everyday devices.
@@ 1. Introduction
Imagine testing whether your drink has been spiked at a bar by simply setting the cup on your tablet screen. Or imagine placing a single drop of water, soda, or milk on your touchscreen to know if it has been contaminated. Finally, imagine detecting trace concentrations of biomolecules such as DNA. While touchscreens have conventionally been designed to reject the effects of liquids such as raindrops, treating it as unwanted noise, this paper takes the opposite approach and investigates if instead commodity capacitive touchscreens can be repurposed to perform liquid sensing. Given the predominance of capacitive touchscreens on modern computing devices, the ability to perform liquid sensing on touchscreens could transform tablets into a portable “lab-on-a-pad” as an accessible means of adulteration and concentration detection in everyday settings, resource-constrained environments and classrooms.
@@
We introduce DropleX, the first system to enable liquid sensing on the integrated capacitive touchscreen of commodity tablets. Our system is capable of detecting microliter-scale liquid samples and performing non-invasive through-container measurements even for sealed liquids. It is able to conduct multiple tests concurrently across the screen, allowing simultaneous characterization of different samples.
However, repurposing the tablet touchscreen to perform liquid sensing is not straightforward and requires tackling several key challenges. First, commercial touchscreens have built-in adaptive filters that continuously recalibrate capacitance readings to a baseline to reject transient effects such as those caused by liquid drops. As a result, any liquid or container placed on the screen is only registered for a few seconds before the readings fades. To address this, we develop a physics-informed mechanism that can disable the filter temporarily for the measurement without any hardware or software modifications.
@@
We present DropleX, the first system that enables liquid sensing on commodity tablet touchscreens without any hardware modifications. We demonstrate adulteration detection accuracies of 96–99% for soda, wine, and milk samples, detection of trace amounts of biochemicals including DNA, NaCl, IPA with accuracies of 93–96%, and through-container sensing with accuracies of 86–96% across multiple containers.
@@
As the proposed approach involves measurements of capacitance heatmaps already accessible to technology manufacturers in software, it can in principle be deployed through a software or firmware update to commodity devices.
@@ 2. Related work
3. Primer on capacitive sensing on touchscreens
Modern multi-point touchscreens (Fig. 2a) are comprised of multiple layers starting from bottom to top:
Touchscreens operate on the idea of mutual capacitance which forms between the two electrode lines (Fig. 2b) (Barrett and Omote, 2010; Hessar et al., 2016). The screen structure is composed of a 2-D array of electrodes from the driving and sensing lines. On the driving line is an AC input drive signal , which creates a shared electric field between the layers, this is sensed by the other layer with the sensing electrodes. We show in Fig. 2c the driving signal of Samsung Galaxy Note 10.1 captured on an oscilloscope (T3DSO1302A, Teledyne LeCroy). The signal is captured by placing the oscilloscope probe directly in contact with the top of the touchscreen and connecting the ground to a metal pole. The driving frequency is approximately 100 kHz with a peak to peak distance is approximately 2.4 mV.
In a touchscreen without any object in contact with it, a mutual capacitive field forms between the driving and sensing electrodes (Fig. 3a). When a conductive object is near, it disturbs the shared electric field. In the case of grounded conductive objects like a human finger that is coupled to Earth ground, the electric field lines are pulled towards the finger and it reduces the capacitance between two electrodes at an intersection (Fig. 3b). The touchscreen controller is able to read the change in capacitance and detect where it has occurred. The human body serves as an effective ground because it is capacitively coupled to Earth, forming a large conductive body with a low-impedance pathway to ground. For the electric field, this provides an easier route to discharge, as the finger is in closer proximity to the screen compared to some of the sensing electrodes which can be farther away and have higher impedance.
In contrast, when placing a liquid sample on the screen it alters the electric field coupling between the driving and sensing electrodes in two ways (Fig. 3c):
(1) where is electrical conductivity of the liquid, is the electrode area, is the separation distance or pitch between electrodes, and is the angular frequency of excitation . The remaining charge is redistributed to neighboring electrodes, effectively reducing the mutual capacitance readings (Zimmerman et al., 1995).
(2) where is vacuum permittivity. Non-conductive or poorly conductive liquids such as oils and alcohols influence the measured capacitance primarily through this effect creating an increase in the measured capacitance.
@@ 4. Enabling liquid sensing on touchscreens
Effect of touchscreen’s liquid rejection filter. While it may appear straightforward to read the capacitive readings from the tablet to perform liquid sensing an immediate challenge becomes apparent when depositing a drop of water on the screen: the capacitive readings fade away. We demonstrate this effect in Fig. 4a which shows the capacitance measured using our tablet’s reported digital units over a period of 25 seconds for two drops of tap water each with a volume of 500 . The capacitance readings are logged at our device’s maximum polling rate of 1.67 Hz corresponding to a frame measured every 0.6 s, implementation details provided in Sec. 7.
@@
The plot reveals two observations, first there is an initial sharp drop in the capacitance readings (using device units). Second, it shows that the measured capacitance reverts back to baseline. For sample 1, it reverts back to baseline quickly by the next frame, while for sample 2, this change was more gradually, taking approximately 18 s.
@@
The reason for this is that manufacturers specifically design software heuristics known as adaptive filters that actively reject liquids such as rain and other spurious events that change the capacitance readings. The ability to disable this adaptive filter from constantly recalibrating to a baseline could enable our system to measure the effect of liquid samples on the capacitance values more reliably.
@@
We observe that the adaptive filter is disabled when a finger, the intended input for touchscreens, is in contact with the display. Specifically, the measured capacitance remains stable over time. This is expected behavior, since the touchscreen must maintain a persistent signal for a valid touch, and recalibrating the finger’s contact back to baseline would cause it to disappear and make normal touch interactions impossible.
@@
Mimicking a permanent touch to disable the filter. It is of course, however, impractical to expect the user to constantly place their finger in contact with the screen throughout the measurement. Our key idea is to mimic a permanent touch event to disable the adaptive filter and enable liquid sensing. Our approach is based on our prior modeling of how a liquid sample affects the touchscreen’s capacitive field (Fig. 3). Specifically, when a liquid is deposited in the screen both its conductivity and permittivity change the capacitive field. When a drop of water is deposited on the screen, the dielectric permittivity term dominates compared to conductance. This is because a sample of liquid is thick, and in the case of water can be 2 mm in height due to its surface tension, which can be thicker than tablet touchscreens which can have a thickness of 1 mm (thi, 2025). Because of this the electric field is largely confined to within the liquid.
@@
If there was a way to modify the liquid so that the conductivity term dominates, it could then mimic the effect of a finger. Specifically, if the liquid height could be reduced to a thin film for the electric field to pass through, the conductivity term would dominate and it would function like a finger. Our key observation is that when drawing up the liquid, a thin film remains which causes the electric field to be shunted to parasitic ground, and is effectively recognized by the screen as a “finger”.
@@
To demonstrate this, we deposit a 500 “priming drop” of water onto an area of the screen which initially induces a positive change in measured capacitance (negative in device units). After this, however upon drawing up the sample using the pipette the capacitance values become negative (positive in device units) and causes a region of pixels to be in a permanently “stuck” position, effectively emulating a permanent touch. The effect on the measured capacitance is illustrated in Fig. 5b where it can be observed that depositing this initial sample results in a sharp drop in capacitance of around -200 device units, and after it is drawn up, it reaches a stable a value of approximately 400 device units.
@@
After this, the adaptive filter is disabled, and the liquid sample of interest can be deposited elsewhere on the screen and its effects can be examined. We show in Fig. 4b that after disabling the effects of the adaptive filter, we deposit two samples of tap water as before. This results in a sharp drop in mutual capacitance, and a transient stabilization period of about one frame, after which the sample readings remain stable across the measurement period of 25 s. We evaluate the stability of these readings over a longer period in Sec. 8.
@@
Importantly, we note that a pipette or specialized applicator is not required to deposit or draw up the priming drop. As show in Fig. 6, the user can simply dip their finger into water, where surface tension forms a small pendant droplet that remains attached to the skin, and gently touch the touchscreen to transfer it onto the surface. The same droplet can then be immediately removed using a regular tissue, which draws the liquid away via capillary wicking as the water is absorbed by its fibers.
@@ 5. Modeling capacitance measurements
Physics-Informed Model. We aim to fit a model that explains these trends, specifically focusing on the data points for the concentrations of IPA and NaCl. To do this, we first leverage a physics-informed model based on the idealized parallel-plate capacitor model Eq. 3. However, this idealized model alone is insufficient for two reasons:
@@ 6.1. Empirical characterization and analysis
6.1.1. Multi-scale spatial sensitivity analysis
To characterize the change in capacitance instead of using the centroid which only measures the value at a single cell, which would make sense if the droplet was deposited only in the center. For drops in other locations which straddle the boundary between multiple cells we sum together the values across cells with a negative change in capacitance. We show in Fig. 11a the magnitude capacitance of the droplet. Both the 10 and 25 samples had comparatively higher capacitance values and lower variation at the center and vertical locations, and lower values at the corner and horizontal locations. The results suggest that placing droplets at the center of the electrode produce the most reliable results. This is the position that we adopt in all experiments. We note however that the use of larger volume samples occupy multiple cells and our system captures features of the spatial information in the region around the centroid of the drop for downstream applications.
with , , and .
The sensitivity map in Fig. 11b, shows two clusters to the left and right of the screen which are sensitive, while the sides and the center have low sensitivity.
@@ 6.1.2. Effect of sample volume
6.1.3. Temporal noise characterization and reduction
Here, we characterize the touchscreen’s sensor noise over time. We note that because the touchscreen is primarily designed to detect touch locations rather than precise capacitance values, it is inherently not optimized for low-noise measurements. To characterize the noise profile, we record the tablet capacitance readings over the course of 30 seconds without anything on the screen. We show in Fig. 13a the standard deviation per pixel across the full screen. We can observe the standard deviation is high in the center of the screen, and lower at the sides, the average standard deviation is 8. To reduce the effect of noise we perform averaging across each frame. Fig. 13b shows the effect of averaging across 20 frames (12 seconds), with an average standard deviation across the screen of 2.8. The relationship between averaging time and the average standard deviation across the screen is shown in Fig. 13c, the plot shows the theoretical relationship of , where is the number of frames being averaged over as well as the measured decrease in standard deviation over 30 seconds. The plot shows that the measured readings decrease with a trend closely following the theoretical model.
@@ 6.1.4. Effect of Interfacial Layers
Through-Container Sensing. Given that the electric field is able to pass through thin films with up to as in the case of PET plastic, we next examine the effects of through-container sensing of liquids in containers (Fig. 14). To do this, we select three different containers: a plastic drinking cup (), a beaker (), and a vial (), and filled them with water until they reached the 2 cm mark and placed them on the touchscreen. The measured heatmaps reveal four points:
@@ 6.2.1. Automatic droplet identification
To automatically identify the location of a drop, we use a -score-based detection algorithm. For each frame, the algorithm computes the mean and standard deviation of all pixel values and applies a negative threshold defined as , where is the mean, is the standard deviation, and is a user-configurable parameter set to 2, reflecting the decreased capacitance effect (in device units) of a droplet on the screen. Cells that fall below this threshold are considered as candidate droplet locations.
Our algorithm then searches for the connected component (cells) around the candidate locations. We consider cells as connected if they share an edge of a corner, including diagonal neighbors (8-connectivity). This search produces a set of candidate regions for the droplet.
To filter out noise, the algorithm removes regions smaller than a minimum size threshold. When considering droplets that are 50 or smaller which only occupy one cell, we set this threshold to 1. The user can set in the app the volume of the drop they wish to deposit, and using prior information about how large of a surface area the droplet could probably spread, the app can set a conservative minimum size threshold.
Finally, regions are further filtered based on geometric constraints. Specifically, we assume that droplets are approximately elliptical in shape and that the ratio between their major and minor axes do not differ significantly. We discard candidate regions if the difference between length and width does exceeds 2 pixels. This approach effectively isolates compact, roughly elliptical regions corresponding to droplet locations.
Instead of running this algorithm continuously, we only do so when a droplet has been detected to have been deposited on the screen. We do this by maintaining a log of frame to frame changes using the metric: , where is the heatmap at timestep . When exceeds a running threshold , we mark that frame as being when a drop has been deposited. We set the sensitivity parameter to be 2.
@@ 6.2.2. Capturing liquid spreading behavior
Prior works (Horstmann et al., 2021) on liquid sensing using capacitive sensing have focused on reporting the capacitance value at the centroid of the liquid deposit location. We note that using this alone is typically insufficient to obtain high classification accuracies at distinguishing between different liquids and classifications, and error bars between different classes overlap. This is due at least in part because a single centroid value does not capture spatial variations in how a liquid spreads.
Our approach instead captures capacitance information in the patch around the centroid which covers a sufficient neighborhood to capture spatial variations caused by different liquid drops while keeping input size computationally manageable. This spatial information implicitly captures the effects of surface tension which affect how the liquid spreads across the surfaces and interacts with neighboring electrodes. For example as shown in Fig. 7, when comparing water () and isopropyl alcohol (), alcohol spreads and flattens more easily. So for the same volume, alcohol occupies a larger surface area and changes the capacitance in a larger neighborhood compared to water.
We designed a compact convolutional neural network to comprehensively extract all these features effectively, specifically designed for single-frame classification. The network includes two convolutional layers with 3×3 kernels, increasing channel depth from 1 to 32 and then 64, each followed by max pooling. Dropout is applied after the convolutional blocks to prevent overfitting. The flattened features pass through a fully connected layer with ReLU activation, another dropout layer, and an output layer for classification. This lightweight model effectively extracts spatial features from individual frames, enabling reliable detection of adulteration types and other liquid categories.
@@ 7.2. Training methodology
We construct a frame-wise training framework in which each time frame from each region is treated as an individual data sample. Compared with region-wise training, which averages all frames within a region and thus loses valuable temporal variation, and sequence-wise training using CNNs combined with LSTM or Transformer architectures, the frame-wise approach makes fuller use of the available data and maximizes the effective dataset size. Moreover, since our measurements contain limited temporal dynamics, employing LSTM or Transformer models to extract temporal features would not provide significant benefits.
Specifically, we extract 50 time frames from each region, thereby expanding the dataset size by a factor of 50. The classifier is trained and evaluated using 5-fold cross-validation to ensure robust performance, with training parameters set to 50 epochs, batch size 16, learning rate 0.001, and a ReduceLROnPlateau scheduler to mitigate overfitting.
@@ 8.1. Adulteration detection
Soda spiking. Here we consider the scenario of a soft drink being spiked. To do this, we use as our base soda Coke, and spike it with different concentrations of industrial ethanol, a form of alcohol unsuitable for human consumption. Ethanol is added at concentrations of 10, 20, 40, 60, and 80% by volume. This scenario simulates intentional contamination with toxic substances. The dataset includes 77 unadulterated and 377 adulterated samples.
Wine adulteration. Next, we investigate the adulteration of wine, which already contains alcohol. Here, the base liquid is red wine, to which industrial ethanol is added in increments of 20%. Compared to soda spiking, this scenario is more challenging because the background alcohol content in wine reduces the relative change caused by adulteration. In total, 88 unadulterated and 231 adulterated samples are collected.
@@
Our results in Fig. 16a show the precision and recall at detecting the adulterated class. We note that these metrics of interest better capture the expected performance compared to accuracy because they emphasize the model’s ability to correctly identify adulterated samples even if it is at the cost of occasional false positives. In our application, the occasional false positive is acceptable as it may simply warrant re-testing, or in the worse case, discarding of the liquid sample. The converse is not acceptable as missing an adulterated product poses a greater risk than a false alarm. Our results show our precision and recall is above 90% for the three adulteration detection tasks.
We show the confusion matrix performance for the three tasks in Fig. 16b–d, the results show that the overall accuracy of these classifiers are all above 96%. In the case of soda spiking the rate of false positives is less than 1%. For wine adulteration detection, we note this is a more challenging case as wine already contains alcohol so when it is contaminated with certain concentrations of alcohol, it is more challenging to detect the adulterated sample, however the overall accuracy is 96.1%. In the case of milk adulteration, the overall classification accuracy reaches 98.1%. Notably, for detergent adulteration, which poses genuine health risks to humans, the false detection rate is 0.0%, indicating that the system can reliably distinguish even subtle forms of contamination. These results collectively demonstrate that our method provides a robust and reliable approach for detecting adulteration in various beverages and liquid foods.

### v4

@@ Abstract.
We present DropleX, the first system that enables liquid sensing using the capacitive touchscreen of commodity tablets. DropleX detects microliter-scale liquid samples, and performs non-invasive, through-container measurements for liquid analysis. These capabilities are made possible by a physics-informed mechanism that disables the touchscreen’s built-in adaptive filters, originally designed to reject the effects of liquid drops such as rain, without any hardware modifications. We model the touchscreen’s sensing capabilities, limits, and non-idealities to inform the design of a signal processing and learning-based pipeline for liquid sensing. Our system achieves 89–99% accuracy in detecting microliter-scale adulteration in soda, wine, and milk, 94–96% accuracy in threshold detection of trace chemical concentrations, and 86–96% accuracy in through-container adulterant detection. These exploratory results demonstrate the potential of repurposing commodity touchscreens as a liquid characterization platform for laboratory settings, food and beverage testing, and chemical analysis applications.
@@ 1. Introduction
Capacitive touchscreens on smart devices such as tablets have conventionally been designed to reject the effects of liquids such as raindrops, treating them as unwanted noise. In this work, we take the opposite approach and investigate whether commodity capacitive touchscreens can instead be repurposed to perform liquid sensing. By analyzing how microliter-scale liquid samples and larger liquid volumes in containers interact with capacitive touchscreens, we explore the feasibility of using these devices for liquid characterization in controlled settings. This approach could enable tablets to serve as accessible liquid analysis platforms for laboratory use, food and beverage quality testing, and chemical characterization in educational or resource-constrained environments.
@@
We introduce DropleX, the first system to enable liquid sensing on the integrated capacitive touchscreen of commodity tablets. Our system is capable of detecting microliter-scale liquid samples and performing non-invasive through-container measurements. It is able to conduct multiple tests concurrently across the screen, allowing simultaneous characterization of different samples.
However, repurposing the tablet touchscreen to perform liquid sensing is not straightforward and requires tackling several key challenges. First, commercial touchscreens have built-in adaptive filters that continuously recalibrate capacitance readings to a baseline to reject transient effects such as those caused by liquid drops. As a result, any liquid or container placed on the screen is only registered for a few seconds before the readings fades. To address this, we develop a physics-informed mechanism that can disable the filter temporarily for the measurement without any hardware modifications.
@@
We present DropleX, the first system that enables liquid sensing on commodity tablet touchscreens without any hardware modifications. Under controlled laboratory conditions, we demonstrate adulteration detection accuracies of 96–99% for soda, wine, and milk samples, detection of trace amounts of biochemicals including DNA, NaCl, ethanol with accuracies of 93–96%, and through-container sensing with accuracies of 86–96% across multiple containers.
@@
As the proposed approach involves measurements of capacitance heatmaps already accessible to technology manufacturers in software, it can in principle be deployed through a software or firmware update to commodity devices, making it suitable as a liquid characterization tool for controlled testing environments.
@@ 3. Enabling liquid sensing on touchscreens
Effect of touchscreen’s liquid rejection filter. While it may appear straightforward to read the capacitive readings from the tablet to perform liquid sensing, an immediate challenge becomes apparent when depositing a drop of water on the screen: the capacitive readings fade away. Fig. 2a shows the capacitance measured using our tablet’s reported digital units over a period of 25 seconds for two drops of 500 water. The capacitance readings are logged at 1.67 Hz corresponding to a frame measured every 0.6 s, implementation details provided in Sec. 6.
@@
The plot reveals two observations, first there is an initial sharp drop in the capacitance readings (in device units). Second, the measured capacitance reverts back to baseline. For sample 1, it reverts back to baseline quickly, while for sample 2, this change was more gradually.
@@
The reason is that manufacturers specifically design software heuristics known as adaptive filters that actively reject liquids such as rain and other spurious events that change the capacitance readings. Disabling this adaptive filter from constantly recalibrating to a baseline would enable our system to measure the effect of liquid samples on the capacitance values more reliably.
@@
We observe that the adaptive filter is disabled when a finger, the intended input for touchscreens, is in contact with the display. This is expected behavior, since the touchscreen must maintain a persistent signal for a valid touch.
@@
Mimicking a permanent touch to disable the filter. It is however impractical to expect the user to constantly place their finger on the screen throughout the measurement. Our key idea is to mimic a permanent touch event to disable the adaptive filter and enable liquid sensing. Our approach is based on our modeling of how a liquid sample affects the touchscreen’s capacitive field (Fig. A.2). Specifically, when a liquid is deposited in the screen both its conductivity and permittivity change the capacitive field. The dielectric permittivity term dominates compared to conductance, because a sample of liquid is thick, and in the case of water can be 2 mm in height due to its surface tension, which can be thicker than tablet touchscreens which can have a thickness of 1 mm (thi, 2025). Because of this the electric field is largely confined to within the liquid.
@@
If the liquid height could be reduced to a thin film for the electric field to pass through, the conductivity term would dominate and it would function like a finger. Our key observation is that when drawing up the liquid, a thin film remains which causes the electric field to be shunted to parasitic ground, and is effectively recognized by the screen as a “finger”.
@@
To demonstrate this, we deposit a 500 “priming drop” of water onto the screen which initially induces a positive change in measured capacitance (negative in device units). After this, upon drawing up the sample using the pipette the capacitance values become negative (positive in device units) and causes a region of pixels to be in a permanently “stuck” position, effectively emulating a permanent touch. Fig. 3b shows that depositing this initial sample results in a sharp drop of around -200 device units, and after it is drawn up, it reaches a stable a value of approximately 400 device units.
@@
After the adaptive filter is disabled, the liquid sample of interest can be deposited elsewhere on the screen without diminishing. We show in Fig. 2b that after disabling the adaptive filter, we deposit two samples of tap water as before. This results in a sharp drop in mutual capacitance, after which the sample readings remain stable.
@@
Importantly, we note that a pipette or specialized applicator is not required to deposit or draw up the priming drop. As show in Fig. 4, the user can simply dip their finger into water, drop the water on the touchscreen, and then immediately remove it using a regular tissue, which draws the liquid away via capillary wicking.
@@ 4. Modeling capacitance measurements
Physics-Informed Model. We aim to fit a model that explains these trends, specifically focusing on the data points for the concentrations of IPA and NaCl. To do this, we first leverage a physics-informed model based on the idealized parallel-plate capacitor model Eq. A.3. However, this idealized model alone is insufficient for two reasons:
@@ 5.1.2. Effect of Interfacial Layers
5.1.3. Through-Container Sensing.
Given that the electric field is able to pass through thin films with up to as in the case of PET plastic, we next examine the effects of through-container sensing of liquids in containers (Fig. 10). To do this, we select three different containers: a plastic drinking cup (), a beaker (), and a vial (), and filled them with water until they reached the 2 cm mark and placed them on the touchscreen. The measured heatmaps reveal four points:
@@ 5.2.1. Automatic droplet identification
To automatically localize liquid droplets, we employ a -score-based detection algorithm. For each frame, pixels with values below a negative threshold (with ) are identified as candidate droplet locations, reflecting the characteristic decrease in capacitance (in device units) caused by liquid contact. Connected components are then extracted using 8-connectivity, and small regions likely due to noise are removed based on a minimum size threshold.
Candidate regions are further filtered based on geometric constraints, retaining only compact regions with approximately elliptical shapes. To reduce unnecessary computation, droplet detection is only triggered when a significant frame-to-frame change is observed, indicating that a droplet has been newly deposited on the screen.
@@ 5.2.2. Capturing liquid spreading behavior
Prior work on capacitive liquid sensing (Horstmann et al., 2021) have focused on reporting the capacitance value at the centroid of the liquid deposit location. We note that using this alone is typically insufficient to distinguish between different liquids and classifications, and error bars between different classes overlap.
Our approach instead extract capacitance measurements from an patch centered at the detected droplet location. This region is large enough to capture spatial spreading behavior while remaining computationally efficient. This spatial information implicitly captures the effects of surface tension which affect how the liquid spreads across the surfaces and interacts with neighboring electrodes. For example as shown in Fig. 5, when comparing water () and isopropyl alcohol (), alcohol spreads and flattens more easily. So for the same volume, alcohol occupies a larger surface area and changes the capacitance in a larger neighborhood compared to water.
We designed a compact CNN to comprehensively extract all these features effectively. The network consists of two convolutional layers with kernels followed by max pooling and dropout. This lightweight architecture effectively extracts discriminative spatial features from individual frames, enabling robust liquid and adulteration classification.
@@ 6.2. Training methodology
We employ drop-wise five-fold cross-validation to prevent data leakage. Data partitioning between train and validation sets are conducted at the drop-level. Specifically, all time frames extracted from a single drop are assigned exclusively to either the training or validation set, never both, with an 80-20 train-test ratio across all samples. This ensures that frames from a single sample do not appear in both splits, and avoids optimistic bias.
To ensure sufficient data for training and validation, we extract 50 time frames per region. The classifier is trained and evaluated using 5-fold cross-validation, with training parameters set to 50 epochs, batch size 16, learning rate 0.001, and a ReduceLROnPlateau scheduler.
@@ 7.1. Adulteration detection
Soda spiking. Here we consider the scenario of a soft drink being spiked. To do this, we use as our base soda Coke, and spike it with 10% volume concentration of industrial ethanol, a form of alcohol unsuitable for human consumption. This scenario simulates intentional contamination with toxic substances. The dataset includes 77 unadulterated and 61 adulterated samples.
Wine adulteration. Next, we investigate the adulteration of wine, which already contains alcohol. Here, the base liquid is red wine, to which industrial ethanol is added at 20% volume concentration. Compared to soda spiking, this scenario is more challenging because the background alcohol content in wine reduces the relative change caused by adulteration. In total, 88 unadulterated and 72 adulterated samples are collected.
@@
Our results in Fig. 12a show the precision and recall at detecting the adulterated class. We note that these metrics of interest better capture the expected performance compared to accuracy because they emphasize the model’s ability to correctly identify adulterated samples even if it is at the cost of occasional false positives. In our application, the occasional false positive is acceptable as it may simply warrant re-testing, or in the worse case, discarding of the liquid sample. The converse is not acceptable as missing an adulterated product poses a greater risk than a false alarm.
We show the confusion matrix performance for the three tasks in Fig. 12b–d, the results show that the overall accuracy of these classifiers are all above 89%. In the case of soda spiking the rate of false positives is less than 2%. For wine adulteration detection, we note this is a more challenging case as wine already contains alcohol so when it is contaminated with certain concentrations of alcohol, it is more challenging to detect the adulterated sample, however the overall accuracy is 89.1%. In the case of milk adulteration, the overall classification accuracy reaches 98.1%. Notably, for detergent adulteration, which poses genuine health risks to humans, the false detection rate is 0.0%, indicating that the system can reliably distinguish even subtle forms of contamination. These results collectively demonstrate that our method provides a robust and reliable approach for detecting adulteration in various beverages and liquid foods.